OpenStudy (anonymous):

I'm not sure how to work this out, are they asking for an inverse? Given the function f(x) = 2x - 5 / 3, which of the below expressions is correct? a. f^-1(x) = 5-2x / 3 b. f^-1(x) = 3x - 5 / 2 c. f^-1(x) = 3x + 5 / 2 d. -2x - 5 / 3

OpenStudy (anonymous):

for d. I forgot the f^-1(x) = at the beginning

OpenStudy (anonymous):

Yes they are asking for an inverse. Do you know how to find it?

OpenStudy (anonymous):

So I would change the f(x) with y and then switch the y and x but I am not sure what to do after that.

OpenStudy (anonymous):

would it be d?

OpenStudy (anonymous):

@NobodyOwens

OpenStudy (anonymous):

Ok so switching it gives: x= (2y-5)/3 multiple both sides by 3 3x = 2y-5 add 5 to both sides 3x+5 = 2y Divide both sides by 2 (3x+5)/2 = y

OpenStudy (anonymous):

Oh my goodness. Awesome! Thank you so much! I was on the verge of figuring it out, I just wasn't sure if I was supposed to divide, add, or multiply first. Thank you!

OpenStudy (anonymous):

Haha no problem.